OpenStudy (anonymous):

solve for u: 3u^2-5u=2

OpenStudy (anonymous):

First make it in standard form : \[3u^2 - 5u - 2 = 0\] Try to make the factors of this equation... Will you???

OpenStudy (anonymous):

after working it out i got u=2 and u=7/3.... im not sure if this is right?

OpenStudy (anonymous):

No, 2 is right but other value of u is not right...

OpenStudy (anonymous):

Let me answer this : Have you done like this: \[3u^2 - 6u + u - 2 = 0\] \[3u(u - 2) + 1(u - 2) = 0\] (u-2)(3u + 1) = 0 Now, Either (u - 2) = 0 or (3u + 1 ) = 0 So, u = 2 or u = -1/3...

OpenStudy (anonymous):

ohh, okay, makes more sense

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Welcome dear...